There were 590 people at play. the admission price was $3.00 for adults and $1.00 for children. The admission receipts were $1070. How many adults and children attended.
Answer by Jstrasner(112)

Hey,
For this one we need to split this problem into two equations:
a (adults) + c (children) = 590
3a + c = 1070
Then we substitute 590 - a for c in the second equation.
3a + 590 - a = 1070
2a + 590 = 1070
2a = 560
a = 280
Then we plug this into the first equation:
280 + c = 590
c = 310
I hope this helps!